# The Minimax Lower Bound of Kernel Stein Discrepancy Estimation

### v1

@@ 3.1 Langevin-Stein KSD on $\mathbb{R}^{d}$
under mild conditions, particularly if the kernel $k$ is $c_{0}$-universal (Carmeli et al., 2010; Sriperumbudur et al., 2010a). In the KSD construction (and throughout the paper when considering the Langevin-Stein KSD), we assume that the kernel $k$ is twice continuously differentiable $k\in\mathcal{C}^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d})$]. Indeed, regarding (5), by the reproducing property for kernel derivatives (Zhou 2008, Theorem 1; Aubin-Frankowski and Szabó 2022, Lemma 1), one can write $(\mathcal{A}_{p_{0}}\mathbf{f})(\mathbf{x})$ as an inner product
@@ 3.2 General KSD
We note that $\Psi_{P_{0}}(x)\in\mathscr{H}$ and $K_{0}(\cdot,x)\in\mathscr{H}_{K_{0}}$ ($x\in\mathcal{X}$) but both yield the same kernel $K_{0}$ [by (25) and (26)].
@@ 4 KSD ESTIMATORS
In this section, we recall existing KSD estimators with established convergence rates alongside their computational complexity. Let $X_{1:n}=(X_{1},\dots,X_{n})$ be an i.i.d. sample from $P$ (shortly, $X_{1:n}\sim P^{n}$) from which $\operatorname{KSD}(P_{0},P)$ is estimated.
@@
V-statistic estimator. Replacing $P$ in (29) with the empirical measure $\hat{P}_{n}=\frac{1}{n}\sum_{j=1}^{n}\delta_{X_{j}}$ yields the V-statistic-based KSD estimator (Chwialkowski et al., 2016) $\widehat{\operatorname{KSD}}^{2}_{V}(P_{0},P)\coloneq\operatorname{KSD}^{2}\!\big(P_{0},\hat{P}_{n}\big)=\frac{1}{n^{2}}\sum_{a,b=1}^{n}K_{0}(X_{a},X_{b})$. This estimator has runtime complexity $\mathcal{O}\!\left(n^{2}\right)$ and under a sub-Gaussian assumption on the Stein feature map, one can show (Kalinke et al., 2025) that it has a convergence rate
@@
The runtime complexity of this estimator is $\mathcal{O}\!\left(mn+m^{3}\right)$. Under a sub-Gaussian assumption on the Stein feature map and appropriate spectral decay of its centered covariance operator, the estimator achieves a convergence rate
@@
The main result of this paper is that no KSD estimator can achieve faster convergence rate than $n^{-1/2}$, specifically showing that the V-statistic and the Nyström-KSD estimators are rate-optimal.
@@ 5 MINIMAX ESTIMATION
However, a good estimator should perform well not just for a single $P_{0}$ and $P$, but uniformly well over a range of plausible distributions. This leads to a worst-case analysis, where one considers the maximum risk of an estimator over a large class of $(P_{0},P)$-pairs. Indeed, we let $\mathcal{T}$ be the set of probability measures such that any $P_{0}\in\mathcal{T}$ satisfies Assumption 1 for a fixed kernel $k$ in the case of Langevin-Stein KSD (resp. satisfies Assumption 2 in the general case), guaranteeing that KSD is well-defined. To each $P_{0}$, we associate the sampling probability measures $\mathcal{S}_{P_{0}}$ for which $\operatorname{KSD}(P_{0},P)$ is finite for any $P\in\mathcal{S}_{P_{0}}$:
@@
and control the r.h.s. using Le Cam’s two-point method (outlined in Theorem 3). In the next section, we establish a positive lower bound on (48) with $s_{n}=\Theta\!\left(n^{-1/2}\right)$, implying lower bounds for the minimax risk of KSD estimation. Further, recalling from Section 4 that known KSD estimation rates are $\mathcal{O}(s_{n})$ with $s_{n}=n^{-1/2}$, our results settle the statistical optimality of these estimator.
@@ 6 RESULTS
Next we present our lower bounds on the minimax estimation of KSD, both for the Langevin-Stein KSD on $\mathbb{R}^{d}$ (Section 6.1) and for general domains (Section 6.2).
@@ 6.1 Langevin-Stein KSD
Let $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a kernel. Assume that $k$ is bounded ($\sup_{\mathbf{x}\in\mathbb{R}^{d}}\sqrt{k(\mathbf{x},\mathbf{x})}<\infty$) and translation-invariant ( $\exists\text{ positive definite }\kappa$ such that $k(\mathbf{x},\mathbf{y})=\kappa(\mathbf{x}-\mathbf{y})$ for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$).
@@
For a Gaussian kernel $k$, our following corollary makes the constant $c>0$ explicit.
@@ 6.2 General KSD
In this section, $(\mathcal{X},\tau_{\mathcal{X}})$ is a topological space and we impose the following additional assumption, ensuring that (i) $(\mathcal{X},\tau_{\mathcal{X}})$ is sufficiently equipped with non-constant continuous bounded functions (used throughout the proof; see Lemma B.4) and that (ii) KSD is valid for at least one $P_{0}$.
@@
[…]
@@
    This result shows that the minimax lower bound of KSD estimation on a general topological space $(\mathcal{X},\tau_{\mathcal{X}})$ is $n^{-1/2}$, given Assumptions 2 and 4; in other words, no KSD estimator can achieve a faster rate.
@@
    We also note that the bound in Theorem 1 is achieved for any $n\in\mathbb{N}_{0}$, while Theorem 2 provides an asymptotic bound for the risk.
@@ 7 PROOF SKETCHES
Both of our results use Le Cam’s two-point method. The core idea of this technique is to reduce the problem of finding a lower bound over a large class of distributions $P_{0}\in\mathcal{T}$ and $P\in\mathcal{S}_{P_{0}}$ to the problem of finding a carefully crafted adversarial sequence of distributions; the key technical challenge and one contribution of our work is the construction of this adversarial sequence. Le Cam’s two-point approach, following directly from Tsybakov (2009a, (2.9) and Theorem 2.2), is as follows.
Theorem 3 (Theorem 2.2; Tsybakov 2009a).
@@ 7.1 Proof Sketch for Theorem 1
by Markov’s inequality, and noticing that all Gaussian distributions are in $\mathcal{T}$ for a bounded $k$, we first obtain the bound $R^{*}_{n}\geq\inf_{\hat{F}_{n}}\sup_{P\in\mathcal{S}_{Q_{0}}}P^{n}\big(\hat{\Delta}_{n}\geq C\big)$, with $Q_{0}=\mathcal{N}\big(\bm{0}_{d},\mathbf{I}_{d}\big)$. Using this probabilistic form, we proceed by applying Le Cam’s method. In our case, this boils down to designing an adversarial distribution pair $(P_{1},P_{2})$ such that
$\big|\operatorname{KSD}(Q_{0},P_{1})-\operatorname{KSD}(Q_{0},P_{2})\big|\geq\frac{2c}{\sqrt{n}}$ (52)
while $\operatorname{KL}\big(P^{n}_{1}\|P^{n}_{2}\big)\leq\alpha$ with $0<c,\alpha<\infty$.
To achieve this goal, we let $P_{1}=\mathcal{N}\big(n^{-1/2}\mathbf{e}_{j},\mathbf{I}_{d}\big)$ and $P_{2}=Q_{0}=\mathcal{N}\big(\bm{0}_{d},\mathbf{I}_{d}\big)$. Notice that, in this case, $\operatorname{KSD}(Q_{0},P_{2})=\operatorname{KSD}(Q_{0},Q_{0})=0$, and thus (52) reduces to $\operatorname{KSD}(Q_{0},P_{1})\geq 2c/\sqrt{n}$.
@@
Therefore, by using that $\operatorname{KSD}(Q_{0},P_{2})=0$ and invoking Lemma 2, we obtain
$\operatorname{KSD}^{2}(Q_{0},P_{1})=n^{-1}\int_{\mathbb{R}^{d}}\left\|{\psi_{P_{1}}(\bm{\omega})}\right\|_{\mathbb{C}}^{2}\mathrm{d}\Lambda(\bm{\omega}),$ (54)
@@
Controlling the KL divergence. Utilizing the known expressions for the KL divergence of product measures and the KL divergence of Gaussians (recalled in Lemma C.1 and Lemma C.2, respectively) yields $\operatorname{KL}(P^{n}_{1}\|P^{n}_{2})\leq 1/2\eqcolon\alpha$ for all $n\in\mathbb{N}_{>0}$.
@@ 7.2 Proof Sketch for Theorem 2
Then, we obtain a lower bound by applying Le Cam’s method with the adversarial distribution pair $P_{1}=P_{0}$ and $P_{2}=P_{n}$. $P_{n}$ is defined as a perturbation of $P_{0}$:
@@
where $\varphi\in\mathcal{C}_{b}(\mathcal{X})\setminus\mathcal{F}_{c}(\mathcal{X})$, $\mathbb{E}[\varphi(X)]=0$, $\varphi\not\equiv 0$, and $\epsilon_{n}=cn^{-1/2}$ with c>0c>0.444The existence of such $\varphi$ is guaranteed by Lemma B.4.
We start by showing that $P_{n}$ belongs to $\mathcal{S}_{P_{0}}$. Then, to apply Le Cam’s method, we establish that $|\operatorname{KSD}(P_{0},P_{1})-\operatorname{KSD}(P_{0},P_{2})|\geq 2cn^{-1/2}$ and that $\operatorname{KL}(P_{1}^{n}\|P_{2}^{n})\leq\alpha$, with $0<\alpha<\infty$.
@@
Therefore, by the formula of the KL divergence of product measures (Lemma C.1), we get the bound $\operatorname{KL}(P_{n}^{n}\|P_{0}^{n})=n\operatorname{KL}(P_{n}\|P_{0})\leq cM<\infty$.
@@
Acknowledgements
JCR is supported by the Ministry of Science, Research, and the Arts Baden-Württemberg, project Algorithm Engineering for the Scalability Challenge (AESC). FK is supported by the pilot program Core-Informatics of the Helmholtz Association (HGF).
@@ A.3 Proof of Theorem 1
$\overset{\text{(b)}}{=}\inf_{\hat{F}_{n}}\sup_{P\in\mathcal{S}_{G_{0}}}P^{n}\!\left(\left|\operatorname{KSD}(Q_{0},P)-\hat{F}_{n}\right|>C\right)\overset{\text{(c)}}{\geq}\inf_{\hat{F}_{n}}\sup_{G\in\mathcal{G}}G^{n}\!\left(\left|\operatorname{KSD}(Q_{0},G)-\hat{F}_{n}\right|>C\right),$ (A.45)
@@
To bring ourselves into the setting of Theorem 3, we let $\mathcal{Y}=\left(\mathbb{R}^{d}\right)^{n}$, $\Theta=\left\{\theta_{\rho}=\operatorname{KSD}(Q_{0},G_{\rho})\,:\,\rho\geq 0\right\}$, d(x,y)=|xy|d(x,y)=|x-y| ($x,y\in\mathbb{R})$, and $\mathcal{P}_{\Theta}=\left\{G^{n}_{\rho}\,:\,\rho\geq 0\right\}=\left\{G^{n}_{\rho}\,:\,G_{\rho}\in\mathcal{G}\right\}=\left\{G^{n}_{\rho}\,:\,\theta_{\rho}\in\Theta\right\}$ therein. Hence, the observed data $X_{1:n}\in\mathcal{Y}$ is distributed as $X_{1:n}\sim G_{\rho}^{n}\in\mathcal{P}_{\Theta}$ for some unknown $\theta_{\rho}\in\Theta$. Let $\hat{F}_{n}=\hat{F}_{n}(X_{1:n})$ be any estimator of $\operatorname{KSD}(Q_{0},G_{\rho})$ based on the $n$ samples $X_{1:n}$.
In this setting, we consider the adversarial pair $(\theta_{\rho},\theta_{0})=\big(\operatorname{KSD}(Q_{0},G_{\rho}),\operatorname{KSD}(Q_{0},G_{0})\big)=\big(\operatorname{KSD}(Q_{0},G_{\rho}),0\big)$ with our choice of $\rho=1/\sqrt{n}$; it remains to lower bound $d(\theta_{\rho},\theta_{0})$ and to upper bound $\operatorname{KL}(G^{n}_{\rho}\|G^{n}_{0})$.
@@
    $d^{2}(\theta_{\rho},\theta_{0})$ $\overset{\text{(a)}}{=}\operatorname{KSD}^{2}(Q_{0},G_{\rho})\overset{\text{(b)}}{=}\rho^{2}\int_{\mathbb{R}^{d}}\left\|{\psi_{G_{\rho}}(\bm{\omega})}\right\|_{\mathbb{C}}^{2}\mathrm{d}\Lambda(\bm{\omega})\overset{\text{(c)}}{=}\rho^{2}\int_{\mathbb{R}^{d}}e^{-\left\|{\bm{\omega}}\right\|_{2}^{2}}\mathrm{d}\Lambda(\bm{\omega})\overset{\text{(d)}}{\geq}\rho^{2}\int_{A}e^{-\left\|{\bm{\omega}}\right\|_{2}^{2}}\mathrm{d}\Lambda(\bm{\omega})$ (A.48)
@@
$\inf_{\hat{F}_{n}}\sup_{G\in\mathcal{G}}G^{n}\!\left(\left|\operatorname{KSD}(Q_{0},G)-\hat{F}_{n}\right|>\frac{c}{\sqrt{n}}\right)\geq\max\left(\frac{e^{-1/2}}{4},\frac{1-\sqrt{1/4}}{2}\right)=\frac{1}{4},$ (A.53)

### v2

@@ 3.1 Langevin-Stein KSD on $\mathbb{R}^{d}$
under mild conditions, particularly if the base kernel $k$ is $c_{0}$-universal (Carmeli et al., 2010; Sriperumbudur et al., 2010a). In the KSD construction (and throughout the paper when considering the Langevin-Stein KSD), we assume that the base kernel $k$ is twice continuously differentiable $k\in\mathcal{C}^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d})$]. Indeed, regarding (5), by the reproducing property for kernel derivatives (Zhou 2008, Theorem 1; Aubin-Frankowski and Szabó 2022, Lemma 1), one can write $(\mathcal{A}_{p_{0}}\mathbf{f})(\mathbf{x})$ as an inner product
@@ 3.2 General KSD
We note that $\Psi_{P_{0}}(x)\in\mathscr{H}$ and $K_{0}(\cdot,x)\in\mathscr{H}_{K_{0}}$ ($x\in\mathcal{X}$) but both yield the same Stein kernel $K_{0}$ [by (25) and (26)].
@@
Remark 1.
KSDs can also be defined using the Petti’s integral (Barp et al., 2024), which, for simplicity, we do not consider in this paper.
@@ 4 KSD ESTIMATORS
In this section, we recall two existing KSD estimators with established convergence rates alongside their computational complexity. Let $X_{1:n}=(X_{1},\dots,X_{n})$ be an i.i.d. sample from $P$ (shortly, $X_{1:n}\sim P^{n}$) from which $\operatorname{KSD}(P_{0},P)$ is estimated.
@@
V-statistic estimator. Replacing $P$ in (29) with the empirical measure $\hat{P}_{n}=\frac{1}{n}\sum_{j=1}^{n}\delta_{X_{j}}$ yields the V-statistic-based KSD estimator (Chwialkowski et al., 2016) $\widehat{\operatorname{KSD}}^{2}_{V}(P_{0},P)\coloneq\operatorname{KSD}^{2}\!\big(P_{0},\hat{P}_{n}\big)=\frac{1}{n^{2}}\sum_{i,j=1}^{n}K_{0}(X_{i},X_{j})$. This estimator has runtime complexity $\mathcal{O}\!\left(n^{2}\right)$ and under a sub-Gaussian assumption on the Stein feature map, one can show (Kalinke et al., 2025) that it has a convergence rate
@@
The runtime complexity of this estimator is $\mathcal{O}\!\left(mn+m^{3}\right)$. Under a sub-Gaussian assumption on the Stein feature map and given an appropriate spectral decay of its centered covariance operator and associated lower bound on $m$, the estimator achieves a convergence rate
@@
The main result of this paper is that no KSD estimator can achieve faster convergence rate than $n^{-1/2}$, specifically showing that the V-statistic and the Nyström-KSD estimators are rate-optimal. [^4] One can also obtain the same $n^{-1/2}$ convergence rate (up to logarithmic factors) in the context of distribution compression with KSD (Li et al., 2024).
@@ 5 MINIMAX ESTIMATION
However, a good estimator should perform well not just for a single $P_{0}$ and $P$, but uniformly well over a range of plausible distributions. This leads to a worst-case analysis, where one considers the maximum risk of an estimator over a large class of $(P_{0},P)$-pairs. Indeed, we let $\mathcal{T}$ be the set of probability measures such that any $P_{0}\in\mathcal{T}$ satisfies Assumption 1 for a fixed base kernel $k$ in the case of Langevin-Stein KSD (resp. satisfies Assumption 2 in the general case), guaranteeing that KSD is well-defined. To each $P_{0}$, we associate the sampling probability measures $\mathcal{S}_{P_{0}}$ for which $\operatorname{KSD}(P_{0},P)$ is finite for any $P\in\mathcal{S}_{P_{0}}$:
@@
and control the r.h.s. using Le Cam’s two-point method (outlined in Theorem 3). In the next section, we establish a positive lower bound on (48) with $s_{n}=\Theta\!\left(n^{-1/2}\right)$, implying lower bounds for the minimax risk of KSD estimation. Further, recalling from Section 4 that known KSD estimation rates are $\mathcal{O}(s_{n})$ with $s_{n}=n^{-1/2}$, our results settle the statistical optimality of these estimators.
@@ 6 RESULTS
Next, we present our lower bounds on the minimax estimation of KSD, both for the Langevin-Stein KSD on $\mathbb{R}^{d}$ (Section 6.1) and for general domains (Section 6.2).
@@ 6.1 Langevin-Stein KSD
Let $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a kernel. Assume that $k$ is translation-invariant ( $\exists\text{ positive definite }\kappa$ such that $k(\mathbf{x},\mathbf{y})=\kappa(\mathbf{x}-\mathbf{y})$ for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$). [^5] Note that translation-invariance implies the boundedness of the kernel ($\sup_{\mathbf{x}\in\mathbb{R}^{d}}\sqrt{k(\mathbf{x},\mathbf{x})}<\infty$).
Remark 2.
Kernels satisfying Assumptions 1 and 3 are, for example, the IMQ kernel and the whole class of twice-differentiable Matérn kernels, in particular, Gaussian kernels.
@@
For a Gaussian base kernel $k$, our following corollary makes the constant $c>0$ explicit.
@@ 6.2 General KSD
In this section, $(\mathcal{X},\tau_{\mathcal{X}})$ is a topological space and we impose the following additional assumption, ensuring that (i) KSD is valid for at least one $P_{0}$ and that (ii) $(\mathcal{X},\tau_{\mathcal{X}})$ is sufficiently equipped with continuous bounded functions (used throughout the proof; see Lemma B.4).
@@
Assume that, for at least one $P_{0}\in\mathcal{T}$, KSD is valid in the sense of (6) for all $P\in\mathcal{S}_{P_{0}}$. In other words, for all $\in\mathcal{S}_{P_{0}}, PP$$\neq P_{0}$ iff. $\operatorname{KSD}(P_{0},P)>0$. Further assume that there exists a $\varphi_{0}\in\mathcal{C}_{b}(\mathcal{X})$ such that there exists no $c\in\mathbb{R}$ such that $\varphi_{0}=c$ holds $P_{0}$-almost surely.
@@
    This result shows that the minimax lower bound of KSD estimation on a general topological space $(\mathcal{X},\tau_{\mathcal{X}})$ is $n^{-1/2}$, given Assumptions 2 and 4; in other words, no KSD estimator can achieve a faster rate in the minimax sense.
@@
    We also note that the bound in Theorem 1 is achieved for any $n\in\mathbb{N}_{>0}$, while Theorem 2 provides an asymptotic bound for the risk.
@@ 7 PROOF SKETCHES
Both of our results use Le Cam’s two-point method. The core idea of this technique is to reduce the problem of finding a lower bound over a large class of distributions $P_{0}\in\mathcal{T}$ and $P\in\mathcal{S}_{P_{0}}$ to the problem of finding a carefully crafted adversarial sequence of distributions; the key technical challenge and one contribution of our work is the construction of this adversarial sequence. Le Cam’s two-point approach, following directly from Tsybakov (2009, (2.9) and Theorem 2.2), is as follows.
Theorem 3 (Theorem 2.2; Tsybakov 2009).
@@ 7.1 Proof Sketch for Theorem 1
by Markov’s inequality, and noticing that all Gaussian distributions are in $\mathcal{T}$ for a bounded $k$, we first obtain the bound $R^{*}_{n}\geq\inf_{\hat{F}_{n}}\sup_{P\in\mathcal{S}_{P_{0}}}P^{n}\big(\hat{\Delta}_{n}\geq C\big)$, with $P_{0}=\mathcal{N}\big(\bm{0}_{d},\mathbf{I}_{d}\big)$ now fixed. Using this probabilistic form, we proceed by applying Le Cam’s method. In our case, this boils down to designing an adversarial distribution pair $(\mathbb{P},\mathbb{Q})$ such that
$\big|\operatorname{KSD}(P_{0},\mathbb{P})-\operatorname{KSD}(P_{0},\mathbb{Q})\big|\geq\frac{2c}{\sqrt{n}}$ (52)
while $\operatorname{KL}\big(\mathbb{Q}^{n}\|\mathbb{P}^{n}\big)\leq\alpha$ with $0<c,\alpha<\infty$.
To achieve this goal, we let $\mathbb{P}=\mathcal{N}\big(n^{-1/2}\mathbf{e}_{j},\mathbf{I}_{d}\big)$ and $\mathbb{Q}=\mathcal{N}\big(\bm{0}_{d},\mathbf{I}_{d}\big)=P_{0}$. Notice that, in this case, $\operatorname{KSD}(P_{0},\mathbb{Q})=\operatorname{KSD}(P_{0},P_{0})=0$, and thus (52) reduces to $\operatorname{KSD}(P_{0},\mathbb{P})\geq 2c/\sqrt{n}$.
@@
Remark 5.
Wynne et al. (2025, Example 4.2) contains a similar expression of the KSD. However, their formula relies on the characteristic functions of the target ($\psi_{P_{0}}$) and the sampling distribution ($\psi_{P}$).
@@
Therefore, by using that $\operatorname{KSD}(P_{0},\mathbb{P})=0$ and invoking Lemma 2, we obtain
$\operatorname{KSD}(P_{0},\mathbb{P})=n^{-1}\int_{\mathbb{R}^{d}}\left\|{\psi_{\mathbb{P}}(\bm{\omega})}\right\|_{\mathbb{C}}^{2}\mathrm{d}\Lambda(\bm{\omega}),$ (54)
@@
Controlling the KL divergence. Utilizing the known expressions for the KL divergence of product measures and the KL divergence of Gaussians (recalled in Lemma C.1 and Lemma C.2, respectively) yields $\operatorname{KL}(\mathbb{Q}^{n}\|\mathbb{P}^{n})\leq 1/2\eqcolon\alpha$ for all $n\in\mathbb{N}_{>0}$.
@@ 7.2 Proof Sketch for Theorem 2
Then, we obtain a lower bound by applying Le Cam’s method with the adversarial distribution pair $\mathbb{P}=P_{0}$ and $\mathbb{Q}=P_{n}$. $P_{n}$ is defined as a perturbation of $P_{0}$: [^6] Similar perturbations are known in the testing literature (Anderson et al., 1994).
@@
where $\varphi\in\mathcal{C}_{b}(\mathcal{X})$, $\mathbb{E}_{P_{0}}[\varphi(X)]=0$, $\varphi$ is not identically zero $P_{0}$-almost surely, and $\epsilon_{n}=cn^{-1/2}$ with c>0c>0.777The existence of such $\varphi$ is guaranteed by Lemma B.4.
We start by showing that $P_{n}$ belongs to $\mathcal{S}_{P_{0}}$. Then, to apply Le Cam’s method, we establish that $|\operatorname{KSD}(P_{0},\mathbb{P})-\operatorname{KSD}(P_{0},\mathbb{Q})|\geq 2cn^{-1/2}$ and that $\operatorname{KL}(\mathbb{P}^{n}\|\mathbb{Q}^{n})\leq\alpha$, with $0<\alpha<\infty$.
@@
Therefore, by the formula of the KL divergence of product measures (Lemma C.1), we get the bound $\operatorname{KL}(\mathbb{Q}^{n},\mathbb{P}^{n})=\operatorname{KL}(P_{n}^{n}\|P_{0}^{n})=n\operatorname{KL}(P_{n}\|P_{0})\leq cM<\infty$.
@@
Acknowledgments
The authors thank Lester Mackey for his constructive comments. FK is supported by the pilot program Core-Informatics of the Helmholtz Association (HGF).
@@ A.3 Proof of Theorem 1
$\overset{\text{(b)}}{=}\inf_{\hat{F}_{n}}\sup_{P\in\mathcal{S}_{G_{0}}}P^{n}\!\left(\left|\operatorname{KSD}(G_{0},P)-\hat{F}_{n}\right|>C\right)\overset{\text{(c)}}{\geq}\inf_{\hat{F}_{n}}\sup_{G\in\mathcal{G}}G^{n}\!\left(\left|\operatorname{KSD}(G_{0},G)-\hat{F}_{n}\right|>C\right),$ (A.45)
@@
To bring ourselves into the setting of Theorem 3, we let $\mathcal{Y}=\left(\mathbb{R}^{d}\right)^{n}$, $\Theta=\left\{\theta_{\rho}=\operatorname{KSD}(G_{0},G_{\rho})\,:\,\rho\geq 0\right\}$, d(x,y)=|xy|d(x,y)=|x-y| ($x,y\in\mathbb{R})$, and $\mathcal{P}_{\Theta}=\left\{G^{n}_{\rho}\,:\,\rho\geq 0\right\}=\left\{G^{n}_{\rho}\,:\,G_{\rho}\in\mathcal{G}\right\}=\left\{G^{n}_{\rho}\,:\,\theta_{\rho}\in\Theta\right\}$ therein. Hence, the observed data $X_{1:n}\in\mathcal{Y}$ is distributed as $X_{1:n}\sim G_{\rho}^{n}\in\mathcal{P}_{\Theta}$ for some unknown $\theta_{\rho}\in\Theta$. Let $\hat{F}_{n}=\hat{F}_{n}(X_{1:n})$ be any estimator of $\operatorname{KSD}(G_{0},G_{\rho})$ based on the $n$ samples $X_{1:n}$.
In this setting, we consider the adversarial pair $(\theta_{\rho},\theta_{0})=\big(\operatorname{KSD}(G_{0},G_{\rho}),\operatorname{KSD}(G_{0},G_{0})\big)=\big(\operatorname{KSD}(G_{0},G_{\rho}),0\big)$ with our choice of $\rho=1/\sqrt{n}$; it remains to lower bound $d(\theta_{\rho},\theta_{0})$ and to upper bound $\operatorname{KL}(G^{n}_{\rho}\|G^{n}_{0})$.
@@
    $d^{2}(\theta_{\rho},\theta_{0})$ $\overset{\text{(a)}}{=}\operatorname{KSD}^{2}(G_{0},G_{\rho})\overset{\text{(b)}}{=}\rho^{2}\int_{\mathbb{R}^{d}}\left\|{\psi_{G_{\rho}}(\bm{\omega})}\right\|_{\mathbb{C}}^{2}\mathrm{d}\Lambda(\bm{\omega})\overset{\text{(c)}}{=}\rho^{2}\int_{\mathbb{R}^{d}}e^{-\left\|{\bm{\omega}}\right\|_{2}^{2}}\mathrm{d}\Lambda(\bm{\omega})\overset{\text{(d)}}{\geq}\rho^{2}\int_{A}e^{-\left\|{\bm{\omega}}\right\|_{2}^{2}}\mathrm{d}\Lambda(\bm{\omega})$ (A.48)
@@
$\inf_{\hat{F}_{n}}\sup_{G\in\mathcal{G}}G^{n}\!\left(\left|\operatorname{KSD}(G_{0},G)-\hat{F}_{n}\right|>\frac{c}{\sqrt{n}}\right)\geq\max\left(\frac{e^{-1/2}}{4},\frac{1-\sqrt{1/4}}{2}\right)=\frac{1}{4},$ (A.53)